OpenStudy (anonymous):

A family of 6 eats at an Italian restaurant. Some order a meatball sub for $9. Some order chicken parmesan for $11. The total bill was $58. How many ordered the meatball sub?

OpenStudy (nurali):

m = nuumber who ordered the meatball sub c = number who order the chicken parmesan m + c = 6 9m + 11c = 58 m + c = 6 c = 6 - m 9m + 11(6 - m) = 58 9m + 66 - 11m = 58 -2m + 66 = 58 -2m = 58 - 66 -2m = -8 m = -8/-2 m = 4